Form the dot product of vector A = 23N 37° and vector B = 14N -35°


1
Expert's answer
2021-03-29T09:01:48-0400

"(\vec A\cdot \vec B)=|\vec A||\vec B|\cos \alpha=23\cdot14\cdot\cos(37°+35°)=99.5\ (N^2)" . Answer